# Large Language Models for Computer-Aided Design: A Survey

### v1

@@ 5.5. Model Evaluation
LLMs can also be used to compare two sources of data. For instance, ChatGPT can be used to estimate similarity between two textual descriptions—one generated by a model and the other being the ground truth. This ability of LLMs is also making its way into the CAD domain. For instance, Du et al. (Du et al., 2024a) developed an evaluation framework, CADBench, which utilized GPT-4o (Hurst et al., 2024) for two complementary evaluation tasks: image-based evaluation, which assessed visual fidelity, and script-based evaluation, which compared objective attributes such as size, color, and material. In (Badagabettu et al., 2024), the visual question-answering score between the user query and the generated isometric image was computed using Clip-FlanT5-XL (Lin et al., 2024). Similarly, in (Tang et al., 2025), Tang et al. fed repaired drawings alongside ground truth drawings into GPT-4 (Achiam et al., 2023) for automatic scoring along three dimensions: legibility, completeness, and tolerance.
Wang et al. (Wang et al., 2025b) leveraged the strong visual understanding capabilities of VLMs to score visual objects, aiding in the construction of preference data, which would otherwise be costly and labor-intensive. The rendered CAD images, along with an instruction detailing the evaluation criteria—divided into three categories: shape quality, shape quantity, and distribution—were passed into the VLM, LLaVA-OV-Qwen2-7B (Li et al., 2024d), to compute the scores.
@@ 5.6. Text Generation
Text generation is one of the most common applications of LLMs. Unsurprisingly, many CAD-related works have also employed LLMs for generating text content. For example, Liu et al. (Liu et al., 2023b) combined DALL-E (Ramesh et al., 2022), GPT-3 (Brown et al., 2020), and CLIP (Radford et al., 2021) within CAD software, enabling users to construct text and image prompts based on their modeling needs. Designers could use text-to-image AI to generate reference images, avoid design bias, prevent design mindset, and stimulate new design considerations. Users would input their design intentions, and GPT-3 would suggest prompts for them to select from. After choosing, the suggestions were rephrased, and DALL-E would generate the final results.
@@
Kodnongbua et al. (Kodnongbua et al., 2023) employed ChatGPT to generate text prompts describing variations of a given model. Similarly, Yuan et al. (Yuan et al., 2024b) introduced a task for semantic commenting of CAD programs. In this task, LLMs segmented programs into code blocks, each corresponding to semantically meaningful parts, and assigned semantic labels to each block. They executed the program to generate a 3D shape and rendered images from ten representative viewpoints. Each image was then translated into a photorealistic image using ControlNet (Zhang et al., 2023). Using ChatGPT, the authors segmented the images into semantically meaningful parts and transformed every bounding box to a pixel-wise segment utilizing segment anything model (Kirillov et al., 2023), linking them to corresponding code blocks. They also explored how ChatGPT could comment on programs based on a given example program, finding that while it performed well at commenting similar programs, it struggled to generalize to new shapes and primitives.
In (You et al., 2024), You et al. employed a VLM to generate semantic part comments for their predicted global base structure. They used GPT-4V (OpenAI, 2023) to interpret an input image and decompose it into semantic parts. Mallis et al. (Mallis et al., 2024) fed multimodal user requests into a VLM, which then outputted a response in the form of a natural language plan. Along similar lines, Rukhovich et al. (Rukhovich et al., 2024) utilized GPT-4o (Hurst et al., 2024) for CAD-specific question answering, allowing users to query the system for information related to CAD models. In another recent effort, Yu et al. (Yu et al., 2024) employed a locally built LLM to analyze users’ requirements and generate initial design ideas. The LLM was used to establish structural requirements and guide how to model the scheme. They also employed finite element analysis to improve data and fine-tuned the LLM to generate easily understandable design improvement suggestions.
@@
An investigation of VLMs for automatically recognizing manufacturing features in CAD design was conducted in (Khan et al., 2024a). The process began by converting a CAD file into three different image views, which, combined with prompts outlining the analysis criteria, served as inputs for five VLMs to identify and evaluate the features. The VLMs included both closed-source models such as GPT-4o (Hurst et al., 2024), Claude-3.5-Sonnet, and Claude-3-Opus (Anthropic, 2024), as well as open-source models like MiniCPM-Llama3-V2.5 (Hu et al., 2024) and Llava-v1.6-mistral-7b (Liu et al., 2024d). A review along conceptually similar lines was also conducted by Sun et al. (Sun et al., 2024), who analyzed key developments in rule-based reasoning (RBR), case-based reasoning (CBR), and large AI models for advancing reusable design in CAD software, summarizing both their advantages and disadvantages. They proposed a hybrid framework combining RBR, CBR and LLMs along with knowledge graphs to enhance reusable CAD design.
@@
Picard et al. (Picard et al., 2023) explored the capabilities of GPT-4V (OpenAI, 2023) and LLaVA 1.6 34B (Liu et al., 2024d) in performing various engineering design tasks, which involved both visual and textual information. These tasks were categorized into four main areas: conceptual design, system-level and detailed design, manufacturing and inspection, and engineering education. Most of these tasks involved generating textual descriptions from images and text prompts. They also assessed GPT-4V’s spatial reasoning abilities, concluding that while the model demonstrated some spatial reasoning capability, it was still limited when compared to human performance. In (Wang et al., 2025c), Wang et al. treated 2D drawings as raster images and employed an image encoder to extract features. They utilized general-purpose language scripts to represent 3D parametric models and leveraged an LLM to autoregressively predict parametric sequences in text form. They developed CAD2Program by fine-tuning an open-source VLM, Mini-InternVL-1.5-2B (Gao et al., 2024), which used InternViT-300M (Chen et al., 2024a) as the vision encoder and InternLM2-1.8B (Cai et al., 2024) as the language model. Relevant applications of LLMs in manufacturing were reviewed by Li et al. in (Li et al., 2024e), evaluating LLM use in various tasks through case studies and examples. While this survey also covered areas outside CAD, it discussed CAD-related tasks including data generation, text-grounded 3D content generation, initial design drafts, and idea generation in aerospace design.
@@ 6. Analysis and Discussion
From the aforementioned state-of-the-art works in Section 5, it can be concluded that nearly all approaches utilize LLMs to generate intermediate representations rather than directly outputting 3D CAD models or 2D CAD drawings. This is likely because directly generating accurate 3D or 2D CAD outputs remains a significant challenge for current models. Common intermediate formats include executable code—such as Python scripts, and parametric data—often structured as JSON files, which can then be parsed or executed to construct the final CAD models.
We also make another key observation that most state-of-the-art methods rely on multimodal inputs, making the use of multimodal LLMs increasingly prevalent. This highlights the potential of different input types for LLM-based automation. Among these input types, text and images are used most frequently. However, other data formats—such as point clouds and sketch sequences—are also explored, reflecting the growing diversity of input modalities being considered in the field.
We also record the frequency with which each LLM is used in the state-of-the-art works that we review. Figure 4 summarizes the number of times different LLMs have been employed in the reviewed contributions. As apparent, GPT-4o is the most frequently used model (11 instances), followed by GPT-4 (8 instances) and GPT-4V (4 instances)—all developed by OpenAI. This trend suggests that, despite being closed-source and potentially requiring payment, LLMs from the GPT family are still widely favored by the research community in CAD domain. A likely reason is their superior performance compared to the other types of LLMs, including many publicly available models.
GPT-4oGPT-4GPT-4VLLaVA-NeXTChatGPTLLaMA-3/3.1VicunaMini-InternVL-1.5-2BInternVL2Mixtral-50BInstructGPTCoCaQwen2.5-Coder-7B-InstructGeminiCode LLaMAGPT-3.5GPT-4 turboBLIP2OpenELMGemmaPhi-2LLaVATinyLLaVAQwen2-1.5BCodeT5+LLaVA-1.5 7BLangGraphClip-FlanT5-XLLLaVA-OV-Qwen2-7BGPT-3Claude 3MiniCPM-Llama3-V2.500222244446666888810101010111111118888444433333333222222221111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111Frequency
Figure 4. Bar chart of LLM usage frequency in CAD-related research (sorted descending).
We also analyze the distribution of reviewed works across different application types. Table 3 summarizes the number of studies attributed to each category. As shown, the majority of contributions focuses on CAD code generation (17 works), followed by text generation (12 works). This suggests that CAD code generation is currently the most prominent and attractive application of LLMs in the CAD domain, and it is expected to continue gaining traction among researchers. Text generation, being a core capability of LLMs, also remains a widely explored area within the research community.
@@
         Application          Works
         Data Generation          7
         CAD Code Generation          17
         Parametric CAD Generation          9
@@
         Text Generation          12
We also analyze the datasets used in each state-of-the-art study. Table 4 presents the datasets employed across the reviewed works. Some datasets are sourced from industry data, while others are synthetically generated by the authors. Many of the industrial datasets are derived from publicly available datasets, such as DeepCAD, Onshape [^4] http://onshape.com, ABC (Koch et al., 2019), and others.

### v2

@@ 5.5. Model Evaluation
LLMs can also be used to compare two sources of data. For instance, ChatGPT can be used to estimate similarity between two textual descriptions: one generated by a model and the other being the ground truth. This ability of LLMs is also making its way into the CAD domain. For instance, Du et al. (Du et al., 2024) developed an evaluation framework, CADBench, which utilized GPT-4o (Hurst et al., 2024) for two complementary evaluation tasks: image-based evaluation, which assessed visual fidelity, and script-based evaluation, which compared objective attributes such as size, color, and material. In (Badagabettu et al., 2024), the visual question-answering score between the user query and the generated isometric image was computed using Clip-FlanT5-XL (Lin et al., 2024). Similarly, in (Tang et al., 2025), Tang et al. fed repaired drawings alongside ground truth drawings into GPT-4 (Achiam et al., 2023) for automatic scoring along three dimensions: legibility, completeness, and tolerance. Wang et al. (Wang et al., 2025a) leveraged the strong visual understanding capabilities of VLMs to score visual objects, aiding in the construction of preference data, which would otherwise be costly and labor-intensive. The rendered CAD images, along with an instruction detailing the evaluation criteria divided into three categories: shape quality, shape quantity, and distribution, were passed into the VLM, LLaVA-OV-Qwen2-7B (Li et al., 2024a), to compute the scores.
@@ 5.6. Text Generation
Text generation is one of the most common applications of LLMs. Unsurprisingly, many CAD-related works have also employed LLMs for generating text content. For example, Liu et al. (Liu et al., 2023b) combined DALL-E (Ramesh et al., 2022), GPT-3 (Brown et al., 2020), and CLIP (Radford et al., 2021) within CAD software, enabling users to construct text and image prompts based on their modeling needs. Designers could use text-to-image AI to generate reference images, avoid design bias, prevent design mindset, and stimulate new design considerations. Users would input their design intentions, and GPT-3 would suggest prompts for them to select from. After choosing, the suggestions were rephrased, and DALL-E would generate the final results. Kodnongbua et al. (Kodnongbua et al., 2023) employed ChatGPT to generate text prompts describing variations of a given model. Similarly, Yuan et al. (Yuan et al., 2024a) introduced a task for semantic commenting of CAD programs. In this task, LLMs segmented programs into code blocks, each corresponding to semantically meaningful parts, and assigned semantic labels to each block. They executed the program to generate a 3D shape and rendered images from ten representative viewpoints. Each image was then translated into a photorealistic image using ControlNet (Zhang et al., 2023a). Using ChatGPT, the authors segmented the images into semantically meaningful parts and transformed every bounding box to a pixel-wise segment utilizing segment anything model (Kirillov et al., 2023), linking them to corresponding code blocks. They also explored how ChatGPT could comment on programs based on a given example program, finding that while it performed well at commenting similar programs, it struggled to generalize to new shapes and primitives.
@@
In (You et al., 2025), You et al. employed a VLM to generate semantic part comments for their predicted global base structure. They used GPT-4V (OpenAI, 2023b) to interpret an input image and decompose it into semantic parts. Mallis et al. (Mallis et al., 2025) fed multimodal user requests into a VLM, which then outputted a response in the form of a natural language plan. They compared three VLMs: GPT-4o (Hurst et al., 2024), GPT-4 mini and GPT-4 Turbo (OpenAI, 2023c). Along similar lines, Rukhovich et al. (Rukhovich et al., 2025) utilized GPT-4o (Hurst et al., 2024) for CAD-specific question answering, allowing users to query the system for information related to CAD models. In another recent effort, Yu et al. (Yu et al., 2024) employed a locally built LLM to analyze users’ requirements and generate initial design ideas. The LLM was used to establish structural requirements and guide how to model the scheme. They also employed finite element analysis to improve data and fine-tuned the LLM to generate easily understandable design improvement suggestions. The system used the generalv3.5 model (iFlytek, 2024) provided by the iFlytek AI platform. An investigation of VLMs for automatically recognizing manufacturing features in CAD design was conducted in (Khan et al., 2024b). The process began by converting a CAD file into three different image views, which, combined with prompts outlining the analysis criteria, served as inputs for five VLMs to identify and evaluate the features. The VLMs included both closed-source models such as GPT-4o (Hurst et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024a), and Claude-3-Opus (Anthropic, 2024b), as well as open-source models like MiniCPM-Llama3-V2.5 (Hu et al., 2024) and Llava-v1.6-mistral-7b (Liu et al., 2024d). A review along conceptually similar lines was also conducted by Sun et al. (Sun et al., 2024), who analyzed key developments in rule-based reasoning (RBR), case-based reasoning (CBR), and large AI models for advancing reusable design in CAD software, summarizing both their advantages and disadvantages. They proposed a hybrid framework combining RBR, CBR and LLMs along with knowledge graphs to enhance reusable CAD design. They mentioned several LLMs, including GPT-4 (Achiam et al., 2023).
@@
Picard et al. (Picard et al., 2025) explored the capabilities of GPT-4V (OpenAI, 2023b) and LLaVA 1.6 34B (Liu et al., 2024d) in performing various engineering design tasks, which involved both visual and textual information. These tasks were categorized into four main areas: conceptual design, system-level and detailed design, manufacturing and inspection, and engineering education. Most of these tasks involved generating textual descriptions from images and text prompts. They also assessed GPT-4V’s spatial reasoning abilities, concluding that while the model demonstrated some spatial reasoning capability, it was still limited when compared to human performance. In (Wang et al., 2025c), Wang et al. treated 2D drawings as raster images and employed an image encoder to extract features. They utilized general-purpose language scripts to represent 3D parametric models and leveraged an LLM to autoregressively predict parametric sequences in text form. They developed CAD2Program by fine-tuning an open-source VLM, Mini-InternVL-1.5-2B (Gao et al., 2024), which used InternViT-300M (Chen et al., 2024a) as the vision encoder and InternLM2-1.8B (Cai et al., 2024) as the language model. Relevant applications of LLMs in manufacturing were reviewed by Li et al. in (Li et al., 2024e), evaluating LLM use in various tasks through case studies and examples. While this survey also covered areas outside CAD, it discussed CAD-related tasks including data generation, text-grounded 3D content generation, initial design drafts, and idea generation in aerospace design. They mentioned GPT-4V (OpenAI, 2023b) and OpenAI’s o1 model (Jaech et al., 2024) as important LLMs for CAD in their paper. The study in (Zhang et al., 2024) examined the impact of ChatGPT-3.5 (OpenAI, 2022) on engineering design education, focusing on concept generation and detailed modeling. It found that while the LLM significantly broadened concept diversity, its textual nature and occasional unreliable outputs limited its effectiveness in detailed CAD modeling, underscoring the irreplaceable value of traditional learning resources and hands-on practice.
Similarly, Garcia et al. (Garcia et al., 2025) leveraged Qwen-2, a pre-trained LLM, for a feature classification task in which ASCII-formatted STL files were treated as language-based data, but the approach failed due to repeated interruptions and out-of-memory errors. Likewise, Yin et al. (Yin et al., 2025) also trained ShapeGPT to perform shape captioning, enabling natural-language descriptions of 3D shapes. In parallel, in (Lee et al., 2026), the authors leveraged LLMs to perform material-layer function assignment and building-object type classification. In the same vein, Doris et al. (Doris et al., 2025b) projected 3D CAD models into 2D through three forms of imagery: multiview CAD renderings, close-up detail views, and engineering drawings. They further introduced DesignQA, a benchmark organized into three components: rule extraction, comprehension, and compliance. They evaluated five LLMs: GPT-4o, GPT-4, Gemini-1.0, Claude-Opus, and LLaVa-1.5. Furthermore, Li et al. (Li et al., 2025b) performed knowledge analysis based on the semantic understanding capabilities of LLMs and generated a set of triples comprising six components: the function set, structure set, constraint set, function–structure relationships, structure–structure relationships, and structure–constraint relationships. Moreover, in (Hu et al., 2025), entities, attributes, and relations are extracted from each CAD model and the corresponding graph structure is computed to obtain a vector representation. Then, Vicuna is employed to parse natural language queries and to perform question answering. In addition, Zhang et al. (Zhang et al., 2025a) captioned $\sim$221k local parts, using vertex-based captioning for simple parts and VLM-based captioning for more complex ones. Authors of (Valente et al., 2025) also evaluated and fine-tuned three VLMs on their visual question answering (VQA) performance, Pixtral-12B (Agrawal et al., 2024), LLaVA, and InternVL, all of which operated on image inputs. In a similar manner, Niu et al. (Niu et al., 2025) used a ViT visual encoder and Qwen2.5 for CAD-based VQA tasks. Similarly, (Man et al., 2025) evaluated a range of multimodal LLMs on spatial reasoning and video understanding tasks derived from CAD designs, including gpt-4.1-2025-04-14, claude-3-7-sonnet-20250219, qwen2.5-vl-72b-instruct (Bai et al., 2025), o3-2025-04-16, and gemini-2.5-pro-preview-05-06. In a like manner, Panta et al. (Panta et al., 2025) analyzed the most recently generated CAD model image after execution, produced a description using GPT-4o, and compared it with the design prompt to evaluate correct model generation.
@@
6. CAD Evaluation
To evaluate the capability of their methods, existing studies have introduced a range of assessment metrics. For instance, You et al. (You et al., 2025) evaluated models using Chamfer Distance, part-segmentation accuracy, and part-segmentation mIoU. Similarly, evaluation in (Rukhovich et al., 2025) also used Chamfer Distance, Intersection over Union, and Invalidity Ratio. Likewise, Mallis et al. (Mallis et al., 2025) employed 2D and 3D Accuracy, and for autoconstraining evaluated performance using Primitive F1 and Constraint F1. They also used Chamfer Distance from (Khan et al., 2024a) and additional metrics from (Xu et al., 2022b), including Coverage, Minimum Matching Distance, and Jensen–Shannon Divergence, along with Invalidity Ratio. They also introduced a VLM-based metric termed the LVM score. Moreover, Xie et al. (Xie and Ju, 2025) used Chamfer Distance (Fan et al., 2017), Invalid Rate, and Gemini 2.0 Flash (Wu et al., 2024c; Zhang et al., 2023b) as evaluation metrics. Following (Rukhovich et al., 2025), (Kolodiazhnyi et al., 2025) also evaluated Chamfer Distance, Intersection over Union, and Invalidity Ratio. In addition, Wang et al. (Wang et al., 2025b) used Chamfer Distance (Fan et al., 2017), Invalidity Ratio, Command Accuracy, and Parameter Accuracy. Furthermore, (Lv and Bao, 2025) evaluated code executability, geometric similarity (Chamfer Distance), and image similarity using DINOv2 (Oquab et al., 2024). Similarly, in (Li et al., 2025a), unconditional CAD generation was evaluated using Coverage, Minimum Matching Distance, Jensen–Shannon Divergence, success ratio $S_{R}$, and Novel score. For text-to-CAD, the authors used reconstruction accuracy $ACC_{T}$ (Li et al., 2024d), computed with command accuracy $ACC_{cmd}$, parameter accuracy $ACC_{param}$ (Wu et al., 2021), and success ratio $S_{R}$. Median Chamfer Distance, Minimum Matching Distance, and Jensen–Shannon Divergence were also employed. CAD captioning was assessed with BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004). CAD deletion used Extract Match, while CAD addition used $ACC_{cmd}$ and $ACC_{param}$. Likewise, (Li et al., 2025d) used Chamfer Distance, Hausdorff Distance, and IoGT, G-Score and proposed a novel metric. Guan et al. (Guan et al., 2025) used mean Chamfer Distance, median Chamfer Distance, and Invalidity Ratio as their metrics.
In (Xu et al., 2025), authors evaluated reconstruction with Chamfer Distance, F-score, and Normal Consistency, topology with Segment Error, Dangling Edge Length, and Self-Intersection Ratio, and assessed enclosure with Flux Enclosure Error. Moreover, (Li et al., 2024d) adopted a self-designed accuracy along with Median Chamfer Distance, Minimum Matching Distance, Jensen–Shannon Divergence, and CT-Score. Furthermore, (Wang et al., 2025a) adopted the F1 score following (Khan et al., 2024a), reporting F1-Sketch and F1-Extrusion. In addition, Liu et al. (Liu et al., 2023b) quantified how frequently participants used LLM-provided prompt suggestions. Similarly, (Wu et al., 2024b) introduces sketch accuracy, entity accuracy, and CAD F1 as metrics. In the same vein, Alrashedy et al. (Alrashedy et al., 2024) used Point Cloud Distance, Hausdorff Distance, and Intersection over the Ground Truth (IoGT), while (Khan et al., 2024b) adopted Feature Name Accuracy, Feature Quantity Accuracy, Hallucination Rate, and Mean Absolute Error. (Zhang et al., 2025b) adopted metrics from (Xu et al., 2022b, 2023), including Coverage, Minimum Matching Distance, and Jensen–Shannon Divergence. Inspired by (Wu et al., 2024c; He et al., 2023), (Khan et al., 2024a) also used F1 scores of primitives and extrusions, Chamfer Distance, and Invalidity Ratio, and additionally performed GPT-4V and user-study evaluations, both measured by accuracy. Likewise, Yin et al. (Yin et al., 2025) used 3D Intersection over Union, Chamfer Distance, F-score@1% (Sokolova et al., 2006), and ULIP similarity (Xue et al., 2023). In (Govindarajan et al., 2026), the authors adopt Invalidity Ratio, F1, Chamfer Distance, Segment Error, Dangling Edge Length, Self-Intersection Ratio, Flux Enclosure Error, and Sphericity Discrepancy, Discrete Mean Curvature Difference, and Extract Euler Characteristic Match. Furthermore, in (Yuan et al., 2025), for text-based CAD editing, the authors evaluated Valid Ratio, Jensen–Shannon Divergence following (Wu et al., 2021; Xu et al., 2022b, 2023), and Chamfer Distance as (Khan et al., 2024a). For evaluating instruction alignment, they adopt Directional CLIP Score from (Gal et al., 2022; Brooks et al., 2023). On the other hand, Yuan et al. (Yuan et al., 2024b) designed a scoring algorithm that checks whether the generated code can be executed and then uses GPT-4o to compare the rendered model view with the target input. In (Yuan et al., 2024a), the authors proposed block accuracy and semantic Intersection over Union (IoU) to evaluate models on commenting CAD programs. Similarly, Badagabttu et al. (Badagabettu et al., 2024) reported success rate, and (Kienle et al., 2025) reported accuracy. Evaluation in (Doris et al., 2025a) employed Valid Syntax Rate and $\mathrm{IOU}_{\mathrm{best}}$. Likewise, Rosnitschek et al. (Rosnitschek et al., 2025) measured code-generation error, validation loss, and training time. In (Li et al., 2024c), Li et al. employed parsing rate and intersection-over-union, while (Li et al., 2025c; Sun et al., 2025) also used the same metrics. However, Garcia et al. (Garcia et al., 2025) reported F1 score, training time, GPU prediction time, and CPU prediction time. In addition, (Doris et al., 2025b) employed F1 score, accuracy, BLEU, ROUGE, and similarity. Likewise, (Madireddy et al., 2025) evaluated processing time, number of correction attempts, a binary status indicator, and overall success rate. In (Jones et al., 2025), authors adopted the average CLIP score as metric. Similarly, Chen et al. (Chen et al., 2025) employed ROUGEL (Lin, 2004), BERTScore (Zhang et al., 2020), and accuracy. Following (Haque et al., 2022), they further used Chamfer Distance, Minimum Matching Distance, Jensen–Shannon Divergence, and Invalidity Ratio. (Kodnongbua et al., 2023) measured precision and recall. In contrast, Picard et al. (Picard et al., 2025) employed several evaluation metrics: self-consistency, transitive violations, accuracy, volume fraction error, floating material error, topology optimization metrics, force magnitude, volume fraction VF, filter radius R, and load application angle. Differently, Li et al. (Li et al., 2025b) took accuracy, time, and valid triplets as their evaluation metrics. However, (Hu et al., 2025) employed query accuracy, question-answering accuracy, and F1-score. On the other hand, (Campos et al., 2025) assessed the models in terms of parameters, duration, number of correct and incorrect answers, and overall accuracy. Conversely, Zhang et al. (Zhang et al., 2025a) evaluated their method using Coverage, Minimum Matching Distance, Jensen–Shannon Divergence, and prediction validity as in (Zhang et al., 2025b), supplemented by vertex-based and VLM-based scores, as well as a realism metric reflecting human judgments. Likewise, Panta et al. (Panta et al., 2025) employed point cloud distance, Hausdorff distance, and IoGT as evaluation criteria. Nevertheless, Sadik et al. (Sadik and Bujny, 2025) introduced a composite complexity score to quantify structural complexity, and a final similarity score to assess structural fidelity. In a different vein, Agarwal et al. (Agrawal et al., 2024) used Average Normalized Levenshtein Similarity (Peer et al., 2024) as the primary evaluation metric, alongside training data usage, word-level accuracy, unit accuracy, and numeric accuracy. In comparison, Wang et al. (Wang et al., 2025c) evaluated model retrieval accuracy, 3D reconstruction quality, and parameter estimation accuracy. For 3D reconstruction, they report precision, recall, and F1 score. For model retrieval, they assess the match percentage, and for parameter estimation, accuracy over correctly retrieved models is reported. Alternatively, (McCarthy et al., 2025) evaluated the proportional improvement. From another standpoint, Du et al. (Du et al., 2024) computed the evaluation metric by averaging the outputs across all criteria. However, Tang et al. (Tang et al., 2025) adopted legibility, completeness, and tolerance, as well as PSNR and SSIM from industrial drawing evaluation. Nonetheless, (Man et al., 2025) employed Command and Parameter Accuracy, Offline Closed-Loop Execution Performance, and Geometric Fidelity.
GPT-4oGPT-4GPT-4VLLaVA-NeXTChatGPTLLaMA-3/3.1VicunaMini-InternVL-1.5-2BInternVL2-26BInternVL2_5-78B-MPOMixtral-50BMistral-7BInstructGPTCoCaQwen2.5-Coder-7B-InstructQwen2.5-3BQwen2.5-VL-72BQwen2-1.5BQwen2-VL-2Blava-onevision-qwen2-7bLLaVA-OV-Qwen2-7BGeminiGemini 2.0Code LLaMAGPT-3.5GPT-3.5 TurboGPT-4 TurboBLIP2OpenELMGemmaGemma3Phi-2LLaVATinyLLaVACodeT5+LLaVA-1.5 7BLangGraphClip-FlanT5-XLGPT-3Claude 3Claude-3.5MiniCPM-Llama3-V2.5GPT-2CodeGPT-smallOpenAI o1GPT-4 miniGPT-4.5Ovis2-34Bgeneralv3.5InternLM2-1.8BDeepSeekGrok 3GPT-4.1OpenAI o302244668810101212141416161616141455332277331122111111111166221133221111552211112222111122111133112233111111553311111122111111111144222211Frequency
Figure 4. Bar chart of LLM usage frequency in CAD-related research.
@@ 7. Analysis and Discussion
As clear from Section 6, commonly adopted metrics for CAD evaluation include accuracy, Chamfer Distance, success rate, Valid/Invalid ratio, Minimum Matching Distance, Jensen–Shannon Divergence, Coverage, F1 score, and Intersection over Union. These metrics cover a range of objectives popular for CAD tasks. Furthermore, from the aforementioned state-of-the-art works in Section 5, it can be concluded that nearly all approaches utilize LLMs to generate intermediate representations rather than directly outputting 3D CAD models or 2D CAD drawings. This is likely because directly generating accurate 3D or 2D CAD outputs remains a significant challenge for current models. Common intermediate formats include executable code, such as Python scripts, and parametric data, often structured as JSON files, which can then be parsed or executed to construct the final CAD models. We also make another key observation that most state-of-the-art methods rely on multimodal inputs, making the use of multimodal LLMs increasingly prevalent. This highlights the potential of different input types for LLM-based automation. Among these input types, text and images are used most frequently. However, other data formats such as point clouds and sketch sequences are also explored, reflecting the growing diversity of input modalities being considered in the field. We also record the frequency with which each LLM is used in the state-of-the-art works that we review. Figure 4 summarizes the number of times different LLMs have been employed in the reviewed contributions. As apparent, GPT-4o and GPT-4 are the most frequently used models (16 and 14 instances respectively), all developed by OpenAI. This trend suggests that, despite being closed-source and potentially requiring payment, LLMs from the GPT family are still widely favored by the research community in CAD domain. A likely reason is their superior performance compared to the other types of LLMs, including many publicly available models. We also analyze the distribution of reviewed works across different application types. Table 3 summarizes the number of studies attributed to each category. As shown, the majority of contributions focuses on CAD code generation (27 works), followed by text generation (24 works) and parametric CAD generation (20 works). This suggests that CAD code generation is currently the most prominent and attractive application of LLMs in the CAD domain, and it is expected to continue gaining traction among researchers. Text generation, being a core capability of LLMs, also remains a widely explored area within the research community. We also analyze the datasets used in each state-of-the-art study. Table 4 presents the datasets employed across the reviewed works. Some datasets are sourced from industry data, while others are synthetically generated by the authors. Many of the industrial datasets are derived from publicly available datasets, such as DeepCAD, Onshape [^3] http://onshape.com, ABC (Koch et al., 2019), and others. Finally, we map each state-of-the-art work to its industrial context. This allows us to identify current active applications of LLMs for CAD, as well as the domains that are relatively underexplored. Table 5 presents the industries most commonly mentioned in the reviewed works. As shown in the table, the manufacturing industry currently attracts the most attention of the research community, whereas the shipbuilding industry has garnered relatively little focus. We also include the textile industry in the table which has no notable associated works. Nevertheless, we predict that future works will be associated to this industry due to the relevance of CAD to textile.
@@
Application Number of Works
Data Generation 14
CAD Code Generation 27
Parametric CAD Generation 20
@@
Text Generation 24
There are also notable limitations faced by the current state-of-the-art techniques. For example, many systems rely on LLMs to interpret the feedback obtained from similarity comparisons between the ground truth and the generated CAD model. In these settings, LLMs can generally determine whether the design is satisfactory, but they often lack the necessary information to understand how it should be improved. To address this gap, some works (Alrashedy et al., 2024; Badagabettu et al., 2024) employed visual question answering pipelines that pose a large number of targeted questions to the model and then synthesize refinement suggestions. Although this strategy provides partial relief, it remains inherently constrained, underscoring the need for more effective and informative feedback mechanisms that can guide LLMs toward actionable geometric corrections in the future.
8. Future Directions